# PHENIX Collaboration

### v1

@@ Acknowledgements.
$N(m_{T})=C{m_{T}}^{\alpha_{m{\rm th}}}\exp\left(-\frac{m_{T}}{T_{\rm eff}}\right)$ (16)
where $C$ is a normalization constant. The polynomial exponent of the thermal spectrum is denoted by $\alpha_{m{\rm th}}$, to distinguish it from the Lévy-exponent $\alpha$. Note that the parameter $\alpha_{m{\rm th}}$ of this paper is the same, as the $\alpha$ parameter of Refs. [34, 40, 41]. The exponent $\alpha_{m{\rm th}}=1-d/2$, where $d$ is the number of spatial dimensions of the expansion; hence its allowed range is $1\leq d\leq 3$. The value of $\alpha_{th}$ is fixed to $\alpha_{m{\rm th}}=0$, which corresponds to an effectively two-dimensional expansion [28, 36]. The same value was used in Ref. [47] by the PHENIX experiment when obtaining $T_{\rm eff}$. Its value is assumed to be independent of centrality and is varied in the systematic studies between $1/2$ and $-1/2$. Also similar to Ref. [47], good quality exponential fits are obtained, with ${mCL}\geq-1.1$%, in each centrality class for charged kaons, protons, and antiprotons in the transverse-mass range of $0.1\leq m_{T}-m\leq 1.0$ GeV. Similar to Ref. [47], the mass dependence of the slopes is also well described with affine linear fits: $T_{\rm eff}(m=T_{f}+m\langle u_{T}\rangle^{2}$, but here these fits have a good CL with ${mCL}\geq 0.1$% in each centrality class.
@@
The kinetic freeze-out temperature is denoted by $T_{f}$. Significant centrality dependent results for the kinetic freeze-out temperature are obtained here. In particular, the value of $T_{f}$ is significantly lower for the 0%–10% centrality class than for other centrality classes. For peripheral centrality classes, the value of $T_{f}$ increases, reaching its upper limit, the value of the chemical freeze-out temperature $T_{m{\rm chem}}$. Fixing $T_{f}$ to a centrality independent constant, results in the above affine linear mass-dependent coefficient fits, similar to Ref. [47], having CLs that are too small, well below the $0.1$% $\leq{m{\rm CL}}$ threshold values. A possible reason for a lower kinetic freeze-out temperatures $T_{f}$ in the 0%–10% centrality class is that a larger volume may cool remarkably longer and to a lower temperature, as noted by Hama and Navarra [65].
$\langle u_{mT}\rangle$
@@
The average radial flow is denoted by $\langle u_{T}\rangle$. This parameter influences the overall slope of the $\lambda(m_{T})$ distribution at higher transverse mass. The centrality dependence of $T_{f}$ and $\langle u_{T}\rangle$ are indicated in Fig 15. Because the pion spectra are distorted by feed-down decays (which are excluded while evaluating $T_{f}$ and $\langle u_{T}\rangle$). However, the fits to the remaining data points gave satisfactory descriptions. Note that the pion halo contains the decay products of long-lived resonances, which include the decay products of $\omega$, $\eta$, $\eta^{\prime}$ and $K_{s}^{0}$ mesons. It is thus important to simulate these decay chains precisely and in agreement with available experimental data. The best values of $T_{f}$ and $\langle u_{T}\rangle$ are obtained by simultaneous fits to the slope parameters of the positively and negatively charged kaons, protons, and antiprotons with the formula $T_{\rm eff}=T_{f}+m\langle u_{T}\rangle^{2}$. These parameters are systematically varied within the uncertainties allowed by these fits, shifting and fixing one of the $(T_{f},\langle u_{T}\rangle)$ pair by one standard deviation and refitting the other parameter to take into account their covariation. The invariant spectra of $K^{\pm}$ and those of protons and antiprotons are simulated using fits with good CL. The mass scaling of these spectra is then utilized. Nearly the same are the mass of the $\eta$ meson compared to charged kaons and the mass of the $\eta^{\prime}$ meson compared to protons and antiprotons. This gives a good basis to describe well the null effect, which is the no in-medium mass modification scenario of the $\eta$ and $\eta^{\prime}$ spectrum. This method has been tested and the test was published in Fig. 11 of Ref. [41]. Extrapolating the mass-scaled simulations to the PHENIX acceptance successfully reproduces the $\eta$ spectrum measured in $\sqrt{s_{{}_{NN}}}=200$ GeV Au$+$Au collisions [66] in the $m_{T}-m_{\eta}\geq 1.75$ GeV range.
$T_{m{\rm cond}}$
@@
The effective temperature of the in-medium $\eta^{\prime}$ condensate is denoted by $T_{m{\rm cond}}$, as in Refs. [34, 40, 41]. This parameter, together with the in-medium mass of the $\eta^{\prime}$ mesons, controls the thermal enhancement of the $\eta^{\prime}$ mesons in the medium. The smaller this number, the smaller the $\eta^{\prime}$ mass drop for a given $\lambda(m_{T})/\lambda_{m{\rm max}}$ measurement. Hence, the most conservative assumption is made for the default value, $T_{m{\rm cond}}=T_{f}$. As this temperature is inside the hot and dense, but already hadronic matter, part of our systematic studies is to vary $T_{m{\rm cond}}$ within the allowed range of $T_{m{\rm chem}}\geq T_{m{\rm cond}}\geq T_{mf}$. As the conservative default value lies at one of the edges of the allowed interval, this choice results in a more than usually asymmetric uncertainty distribution on the physical fit parameters, such as the in-medium mass of the $\eta^{\prime}$ and the slope parameter of the $\eta^{\prime}$ after the decay of the condensate, $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$. From these systematic uncertainty studies, the in-medium $\eta^{\prime}$ mass is observed to be smaller than our quoted best values, but cannot easily be larger than the best values because its upper uncertainties are much smaller than its lower uncertainties.
@@
Figure 15: Centrality dependent expectations for $\lambda(m_{T})/\lambda_{m{\rm max}}$, based on Monte-Carlo simulations without any in-medium $\eta^{\prime}$ modification.
@@
For each of the parameters $\alpha_{m{\rm th}}$, $T_{f}$, $\langle u_{T}\rangle$, and $T_{m{\rm cond}}$, the $\chi^{2}$ scans and the corresponding CL maps are redone for the variations. The relative difference between the default and the alternative setting is calculated. The final systematic uncertainty is the quadratic sum of these differences.
@@
Additionally, several other cases are also investigated, including two special cases, $T_{f}=140$ MeV fixed, independent of centrality, and $T_{f}=177$ MeV fixed, also independent of centrality. The first case corresponds to Landau’s calculation of the freeze-out temperature (equal to the pion mass, mass of the lightest neutral quanta), The second case is consistent with the PHENIX publication on charged pion, kaon, and (anti)proton spectra [47]. The latter choice is found to be inconsistent with $T_{m{\rm cond}}\geq T_{f}$. The former gives results that are within the quoted systematic uncertainties. Therefore, these two cases are not included as additional sources of systematic uncertainty.
@@
The simulation does not include the PHENIX detector system; hence, the experimental systematic uncertainties are accounted for by propagating the total systematic uncertainties of the measured $\lambda/\lambda_{m{\rm max}}$. The values of the $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$ parameters with statistical and systematic uncertainties are shown in Fig. 12 and Fig. 13. The centrality-dependent results for the in-medium $\eta^{\prime}$ masses are shown on Fig. 13 and the corresponding numerical values are tabulated in Table 2.
@@
Table 2: The centrality-dependent, fitted values, in [MeV], of the in-medium $\eta^{\prime}$ mass $m^{*}_{\eta^{\prime}}$ and the inverse slope $B^{-1}_{\eta^{\prime}}$. The statistical (stat) and systematic (syst) uncertainties are also shown.
@@
As mentioned in Section VI.1, the normalized intercept parameter dependence on $m_{T}$ and centrality in are investigated in detail. An explanation of the dependence is given in terms of radial flow and $p_{T}$ sharing among the pions arising in $\eta^{\prime}$ decay. The simulations suggest a $\lambda(m_{T})/\lambda_{m{\rm max}}$ curve, that changes monotonically with centrality. These changes appear related to the monotonic decrease of radial flow $\langle u_{T}\rangle$, coupled to a monotonic increase of the kinetic freeze-out temperature $T_{f}$ as the collisions change from most central to more and more peripheral collisions [34]. The values of $T_{f}$ and $\langle u_{T}\rangle$ are obtained from affine linear fits with $T_{\rm eff}=T_{f}+m\langle u_{T}\rangle^{2}$ to the slope parameters of the charged kaon, $K^{+}$ and $K^{-}$, as well as to the proton and the antiproton single-particle spectra. Particular attention is paid to the requirement that the single exponential fit to the single-particle spectra have acceptable CLs with ${m{\rm CL}}\geq 0.1\%$, and that the affine linear fits with $T_{\rm eff}=T_{f}+m\langle u_{T}\rangle^{2}$ also reflect the slopes of the single-particle spectra $T_{\rm eff}$ with a CL $\geq 0.1$%.
@@
The method of propagation of statistical and systematic uncertainties was cross checked though comparison with the less direct methods of Refs. [40, 41] as well as with the PHENIX method of Ref. [67] and gives results that are consistent (and typically smaller) with those presented in the body of this manuscript. Hence the central values, with statistical and systematic uncertainties, presented in the manuscript are obtained with the most conservative of the four different methods that were tested.

### v2

@@ Acknowledgements.
$N(m_{T})=C{m_{T}}^{\alpha_{{\rm th}}}\exp\left(-\frac{m_{T}}{T_{\rm eff}}\right)$ (16)
where $C$ is a normalization constant. The polynomial exponent of the thermal spectrum is denoted by $\alpha_{{\rm th}}$, to distinguish it from the Lévy-exponent $\alpha$. Note that the parameter $\alpha_{{\rm th}}$ of this paper is the same, as the $\alpha$ parameter of Refs. [36, 43, 44]. The exponent $\alpha_{{\rm th}}=1-d/2$, where $d$ is the number of spatial dimensions of the expansion; hence its allowed range is $1\leq d\leq 3$. The value of $\alpha_{th}$ is fixed to $\alpha_{{\rm th}}=0$, which corresponds to an effectively two-dimensional expansion [30, 39]. The same value was used in Ref. [50] by the PHENIX experiment when obtaining $T_{\rm eff}$. Its value is assumed to be independent of centrality and is varied in the systematic studies between $1/2$ and $-1/2$. Also similar to Ref. [50], good quality exponential fits are obtained, with CL $\geq 0.1$%, in each centrality class for charged kaons, protons, and antiprotons in the transverse-mass range of $0.1\leq m_{T}-m\leq{1.0}$ GeV. Similar to Ref. [50], the mass dependence of the slopes is also well described with affine linear fits: $T_{\rm eff}(m)=T_{f}+m\langle u_{T}\rangle^{2}$, but here these fits have a good CL with CL $\geq 0.1$% in each centrality class.
@@
The kinetic freeze-out temperature is denoted by $T_{f}$. Significant centrality dependent results for the kinetic freeze-out temperature are obtained here. In particular, the value of $T_{f}$ is significantly lower for the 0%–10% centrality class than for other centrality classes. For peripheral centrality classes, the value of $T_{f}$ increases, reaching its upper limit, the value of the chemical freeze-out temperature $T_{{\rm chem}}$. Our results for the variations of $T_{f}$ and $\langle u_{T}\rangle$ with centrality are shown as inserted values in Fig. 15. Fixing $T_{f}$ to a centrality-independent constant, results in the above affine linear-mass-dependent coefficient fits, similar to Ref. [50], having CLs that are too small, well below the $0.1$% $\leq{{\rm CL}}$ threshold values. A possible reason for a lower kinetic freeze-out temperatures $T_{f}$ in the 0%–10% centrality class is that a larger volume may cool remarkably longer and to a lower temperature, as noted by Hama and Navarra [67].
$\langle u_{T}\rangle$
@@
The average radial flow is denoted by $\langle u_{T}\rangle$. This parameter influences the overall slope of the $\lambda(m_{T})$ distribution at higher transverse mass and results in a centrality dependent expectation for $\lambda(m_{T})/\lambda_{\textmd{max}}$. Note that the pion halo contains the decay products of long-lived resonances, which include the decay products of $\omega$, $\eta$, $\eta^{\prime}$ and $K_{s}^{0}$ mesons. It is thus important to simulate these decay chains precisely and in agreement with available experimental data. The best values of $T_{f}$ and $\langle u_{T}\rangle$ are obtained by simultaneous fits to the slope parameters of the positively and negatively charged kaons, protons, and antiprotons with the formula $T_{\rm eff}=T_{f}+m\langle u_{T}\rangle^{2}$ where $m$ is the mass of the meson or baryon. These parameters are systematically varied within the uncertainties allowed by these fits, shifting and fixing one of the $(T_{f},\langle u_{T}\rangle)$ pair by one standard deviation and refitting the other parameter to take into account their covariation. The invariant spectra of $K^{\pm}$ and those of protons and antiprotons are simulated using fits with good CL. The mass scaling of these spectra is then utilized. Nearly the same are the mass of the $\eta$ meson compared to charged kaons and the mass of the $\eta^{\prime}$ meson compared to protons and antiprotons. This gives a good basis to describe well the null effect, which is the no in-medium mass modification scenario of the $\eta$ and $\eta^{\prime}$ spectrum. This method has been tested and the test was published in Fig. 11 of Ref. [44]. Extrapolating the mass-scaled simulations to the PHENIX acceptance successfully reproduces the $\eta$ spectrum measured in $\sqrt{s_{{}_{NN}}}=200$ GeV Au$+$Au collisions [68] in the $m_{T}-m_{\eta}\geq 1.75$ GeV range.
$T_{\rm cond}$
@@
The effective temperature of the in-medium $\eta^{\prime}$ condensate is denoted by $T_{{\rm cond}}$, as in Refs. [36, 43, 44]. This parameter, together with the in-medium mass of the $\eta^{\prime}$ mesons, controls the thermal enhancement of the $\eta^{\prime}$ mesons in the medium. The smaller this number, the smaller the $\eta^{\prime}$ mass drop for a given $\lambda(m_{T})/\lambda_{\rm max}$ measurement. Hence, the most conservative assumption is made for the default value, $T_{\rm cond}=T_{f}$. As this temperature is inside the hot and dense, but already hadronic matter, part of our systematic studies is to vary $T_{\rm cond}$ within the allowed range of $T_{\rm chem}\geq T_{\rm cond}\geq T_{\rm f}$. As the conservative default value lies at one of the edges of the allowed interval, this choice results in a more than usually asymmetric uncertainty distribution on the physical fit parameters, such as the in-medium mass of the $\eta^{\prime}$ and the slope parameter of the $\eta^{\prime}$ after the decay of the condensate, $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$. From these systematic uncertainty studies, the in-medium $\eta^{\prime}$ mass is frequently observed to be smaller than our quoted best values, but cannot easily be larger than the best values because its upper uncertainties are much smaller than its lower uncertainties.
@@
Figure 15: Centrality dependent expectations for $\lambda(m_{T})/\lambda_{{\rm max}}$, based on Monte-Carlo simulations without any in-medium $\eta^{\prime}$ modification.
@@
For each of the parameters $\alpha_{{\rm th}}$, $T_{f}$, $\langle u_{T}\rangle$, and $T_{{\rm cond}}$, the $\chi^{2}$ scans and the corresponding CL maps are redone for the variations. The relative difference between the default and the alternative setting is calculated. The final systematic uncertainty is the quadratic sum of these differences.
@@
Additionally, several other cases are also investigated, including two special cases, $T_{f}=140$ MeV fixed, independent of centrality, and $T_{f}=177$ MeV fixed, also independent of centrality. The first case corresponds to Landau’s calculation of the freeze-out temperature (equal to the pion mass, mass of the lightest neutral quanta). The second case is consistent with the PHENIX publication on charged pion, kaon, and (anti)proton spectra [50]. The latter choice is found to be inconsistent with $T_{{\rm cond}}\geq T_{f}$. The former gives results that are within the quoted systematic uncertainties.
Similarly to Ref. [1], the statistical uncertainty of $\lambda_{\rm max}$ is treated as a normalization uncertainty. Both this uncertainty ($\approx{1}$%) and the systematic uncertainty ($\approx{1}$%) caused by the choice of $m_{T}$ range when calculating $\lambda_{\rm max}$ are negligible compared to other uncertainties, except in the 50%–60% centrality case of $B_{\eta^{\prime}}^{-1}$. In that case, no significant in-medium mass modification is found. Thus, $B_{\eta^{\prime}}^{-1}$ cancels from the results and cannot be precisely determined. Therefore, these three negligibly small sources of systematic errors are not included here.
@@
The simulation does not include the PHENIX detector system; hence, the experimental systematic uncertainties are accounted for by propagating the total systematic uncertainties of the measured $\lambda/\lambda_{{\rm max}}$. The values of the $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$ parameters with statistical and systematic uncertainties are shown in Fig. 13 and Fig. 12, respectively. The centrality-dependent results for the in-medium $\eta^{\prime}$ masses are shown on Fig. 13 and the corresponding numerical values are tabulated in Table 2.
@@
Table 2: The centrality-dependent, fitted values, in [MeV], of the in-medium $\eta^{\prime}$ mass $m^{*}_{\eta^{\prime}}$ and the inverse slope $B^{-1}_{\eta^{\prime}}$. The statistical (stat) and systematic (syst) uncertainties are also presented.
@@
As mentioned in Section VI.1, the normalized intercept parameter dependence on $m_{T}$ and centrality in are investigated in detail. An explanation of the dependence is given in terms of radial flow and $p_{T}$ sharing among the pions arising in $\eta^{\prime}$ decay. The simulations suggest a $\lambda(m_{T})/\lambda_{{\rm max}}$ curve, that changes monotonically with centrality. These changes appear related to the monotonic decrease of radial flow $\langle u_{T}\rangle$, coupled to a monotonic increase of the kinetic freeze-out temperature $T_{f}$ as the collisions change from most central to more and more-peripheral collisions [36]. The values of $T_{f}$ and $\langle u_{T}\rangle$ are obtained from affine linear fits with $T_{\rm eff}=T_{f}+m\langle u_{T}\rangle^{2}$ to the slope parameters of the charged kaon, $K^{+}$ and $K^{-}$, as well as to the proton and the antiproton single-particle spectra. Particular attention is paid to the requirement that the single exponential fit to the single-particle spectra have acceptable CLs with ${{\rm CL}}\geq 0.1\%$, and that the affine linear fits with $T_{\rm eff}=T_{f}+m\langle u_{T}\rangle^{2}$ also reflect the slopes of the single-particle spectra $T_{\rm eff}$ with a CL $\geq 0.1$%.
@@
The method of propagation of statistical and systematic uncertainties was cross checked though comparison with the less direct methods of Refs. [43, 44] as well as with the PHENIX method of Ref. [69]. These methods gave results that are consistent with (and typically have smaller uncertainties than) those presented in the body of this manuscript. Hence the central values, with statistical and systematic uncertainties, presented in the manuscript are obtained with the most conservative of the four different methods that were tested.